the number of crests in a light wave traveling from the sun to the Earth.

The distance from Earth to the sun is 1.5*10^8 km.
Find the number of crests in a light wave of frequency 1.4×1014 s^-1 traveling from the sun to the Earth.
Answer #1
given v=f?, v=c=3x10^8 m/s

?=v/f

?=c/f

?=3x10^8 / 1.4x10^14

?=2.14285714x10^-6

So

1.5x10^8 / 2.14285714x10^-6=7x10^13 waves
answered by: Autumn Winchester